OpenStudy (anonymous):

An electronics store makes a profit of $72 for every television sold and $90 for every computer sold. The manager’s target is to make at least $360 a day on sales from televisions and computers. Write a linear inequality and graph the solutions. What are three possible solutions to the problem?

OpenStudy (agent47):

72x+90y>=360

OpenStudy (agent47):

8x+10y>=40 4x+5y>=20

OpenStudy (anonymous):

@saifoo.khan Can you help with this graph or do you not understand it either

OpenStudy (saifoo.khan):

What do you don't understand? @chicagochica5

OpenStudy (anonymous):

I dont know how to figure it out @saifoo.khan

OpenStudy (saifoo.khan):

Like @Agent47 said: Suppose TVs as "x" and computers as "y".

OpenStudy (saifoo.khan):

72x+90y>=360

OpenStudy (saifoo.khan):

72 TVs minimum should be sold and 90 computers..

OpenStudy (anonymous):

I get that part

OpenStudy (saifoo.khan):

Which will help him gain a profit of $360 daily..

OpenStudy (saifoo.khan):

Okay.. then..?

OpenStudy (anonymous):

I dont know if it should be a dotted or staight line

OpenStudy (saifoo.khan):

<= or >= should be solid. < or > should be dotted

OpenStudy (anonymous):

Would the answer be C?

OpenStudy (anonymous):

Would it be C @saifoo.khan

OpenStudy (saifoo.khan):

Not c. You are doing the shading part incorrect.

OpenStudy (anonymous):

Oh, Sorry I really do not understand what Im doing. So would it be A then?

OpenStudy (anonymous):

@saifoo.khan would it be A then?

OpenStudy (anonymous):

Actually would it be D @saifoo.khan

OpenStudy (anonymous):

@Agent47 Would the answer be D?

OpenStudy (anonymous):

Omg, I am not understanding any of this

OpenStudy (agent47):

hang on, let's take a look at this.. 4x+5y>=20 5y>=20-4x y>=4-(4/5)x

OpenStudy (agent47):

the dashed lines are automatically discarded, since you have more than or equal to.

OpenStudy (agent47):

When x = 0, y>=4, so the y intercept is 4

OpenStudy (anonymous):

Oh, So it would be A like I said 22 minutes ago

OpenStudy (agent47):

Yea, I'm sorry, i was afk

OpenStudy (anonymous):

Thanks @Agent47